# Weak interaction axial form factors of the octet baryons in nuclear medium

### v1

@@ V.2.2 Results for $G_{P}^{\ast}$
The results of our calculations for the induced pseudoscalar form factor $G_{P}$ in the nuclear medium are presented on the each right side of Fig. 6 for $\Delta I=1$ and Figs. 7 and 8 for $\Delta S=1$. The ratios $G_{P}^{\ast}/G_{P}$ are presented on the each left side of Figs. 11 and 12. The range of $Q^{2}$ is adjusted to each case in order to better visualize the shape of the functions.
@@
On the each right side of Figs. 6, 7 and 8, we can still observe the enhancement of $G_{P}$ near $Q^{2}=0$ as in vacuum, for both $\Delta I=1$ and $\Delta S=1$. Since, as mentioned already, the $G_{P}$ estimates at $Q^{2}=0$ may be inaccurate, we avoid definitive conclusions about the point $Q^{2}=0$, and for the ratios $G_{P}^{\ast}/G_{P}$ in Figs. 11 and 12, we display only the region $Q^{2}>0.012$ GeV2.
@@
In general, one can observe an overall suppression of the form factors $G_{P}^{\ast}$ for both cases ($\Delta I=1$ and $\Delta S=1$) compared with the vacuum, particularly at low $Q^{2}$. The suppression increases with the density. The suppression at low $Q^{2}$ is diluted when $Q^{2}$ increases. In some cases, there are apparent overlaps of the different calculations ($\rho=0$, 0.5 $\rho_{0}$ and $\rho_{0}$), particularly for heavy baryons. See last two lines at the right side of Fig. 6 and all lines at the right side of Fig. 8. The apparent overlaps are the consequence the reduced scale in $Q^{2}$ and the thickness of the lines.
@@
The analysis of the suppression effect in medium on $G_{P}$ is simplified when we look at the ratios $G_{P}^{\ast}/G_{P}$ that are presented on the left side of Fig. 11 for $\Delta I=1$ and on the left side of Fig. 12 for $\Delta S=1$.
@@
In some cases ($\Sigma^{+}\to\Lambda$, $\Sigma^{0}\to\Sigma^{+}$, $\Xi^{-}\to\Lambda$ and $\Xi^{-}\to\Sigma^{+}$) we have more significant reductions of $G_{P}^{\ast}/G_{P}$ at low $Q^{2}$ due to the more effective reduction of the meson cloud component, as discussed already for $G_{A}^{\ast}$ (recall that $G_{P}^{\ast}\propto G_{A}^{\ast}$). At large $Q^{2}$, we can expect for sufficient large values of $Q^{2}$ the dominance of the pole term (proportional to $G_{A}^{{\rm B}\ast}$). This effect can be observed in the figures for the vacuum (each left side of Figs. 6, 7 and 8) in the upper limit of $Q^{2}$.
@@
We start the discussion with the $\Delta I=1$ transitions. For the transitions $\Delta I=1$, on the left side of Fig. 11, the significant reduction of the ratio $G_{P}^{\ast}/G_{P}$ can be understood as a consequence of the dominance of the pole and meson cloud terms, a sum proportional to $G_{A}^{\ast}=G_{A}^{{\rm B}\ast}+G_{A}^{{\rm MC}\ast}$. In that case $G_{P}^{\ast}\propto(M_{BB^{\prime}}^{\ast})^{2}G_{A}^{\ast}$, since in the present model $m_{\pi}$ has the same value for the vacuum and medium ($m_{\pi}$ is almost insensitive to the medium modifications [17, 18, 19, 20]).
@@
The result $G_{P}^{\ast}/G_{P}<1$ is then naturally explained by the mass reduction in medium ($\frac{M_{B}^{\ast}+M_{B^{\prime}}^{\ast}}{M_{B}+M_{B^{\prime}}}<1$) and by the reduction of $G_{A}$ in medium ($G_{A}^{\ast}/G_{A}<1$) discussed already. The suppression is more effective due to the power two in the mass ratio. One can then expect a triple suppression of the ratio $G_{P}^{\ast}/G_{P}$ in medium for $\Delta I=1$. The impact of the suppression is smaller for heavy baryons (mass ratio and $G_{A}^{\ast}/G_{A}$ are reduced more weakly), and lower densities $\rho$. Note, in this respect that the reduction of $G_{P}^{\ast}/G_{P}$ is very mild for the $\Xi^{-}\to\Xi^{0}$ near the upper limit of $Q^{2}$.
We consider now the $\Delta S=1$ transitions. In this case, the pole and the meson cloud term include the factor $\frac{(M_{B}^{\ast}+M_{B^{\prime}}^{\ast})^{2}}{m_{K}^{*2}+Q^{2}}$. At low $Q^{2}$ the effect of mass reduction of the kaon cancels in part of the effect of the reduction of the baryon masses in medium (factor $\frac{(M_{B}^{\ast}+M_{B^{\prime}}^{\ast})^{2}}{(M_{B}+M_{B^{\prime}})^{2}}$), since $\frac{m_{K}^{2}}{m_{K}^{\ast 2}}>1$. Thus, at low $Q^{2}$, the ratio is dominated by the reduction of the meson cloud contribution (more significant reduction at low $Q^{2}$). For large $Q^{2}$, the effect of the kaon mass modification $m_{K}\to m_{K}^{\ast}$ is included in the factor $1+\frac{m_{K}^{2}-m_{K}^{*2}}{m_{K}^{*2}+Q^{2}}$ and decreases quickly to the unit. The consequence is that ratio $G_{P}^{\ast}/G_{P}$ can be described by Eq. (38), with $G_{A}^{\ast}$ replaced by $G_{A}^{B\ast}$ (dominance of the pole term at large $Q^{2}$). The conclusion is then that at large $Q^{2}$, we can expect a stronger suppression for light baryons (more significant baryon mass effect and reduction of $G_{A}^{{\rm B}\ast}/G_{A}^{\rm B}$) and almost no medium effects for heavy baryons ($G_{P}^{\ast}/G_{P}\approx 1$).
@@ V.3.1 Axial-vector form factor $G_{A}^{\ast}$
The results of the calculation of $G_{A}^{\ast}(Q^{2})/G_{A}^{\ast}(0)$ are presented on the each right side of Figs. 9 and 10, for $\Delta I=1$ and $\Delta S=1$, respectively.
@@
As mentioned already, the slope of $G_{A}^{\ast}(Q^{2})/G_{A}^{\ast}(0)$ is related to the axial-vector square radius $r_{A}^{2}\right>$ in the limit $Q^{2}=0$. The results displayed on Figs. 9 and 10, suggest, however, that slope is smaller than that we can expect from a typical dipole form (2) with $M_{A}\simeq 1.05$ GeV: $r_{A}^{2}\right>=\frac{12}{M_{A}^{2}}\simeq 0.42$ fm2 for the $n\to p$ transition. The axial-vector square radius obtained for the vacuum are of the order of 0.1 fm2 for the $n\to p$, $\Sigma^{0}\to\Sigma^{+}$ transitions; 0.16 fm2 for $\Xi^{-}\to\Sigma^{0}$ and $\Xi^{0}\to\Sigma^{+}$ transitions; 0.2 fm2 for the $\Sigma^{+}\to\Lambda$ and about 0.35 fm2 for the $\Xi^{-}\to\Xi^{0}$, $\Lambda\to p$, $\Sigma^{-}\to n$ and $\Sigma^{0}\to p$. The $\Xi^{-}\to\Lambda$ transition is more atypical and has a negative axial-vector square radius (as can be inferred from the figure). These results are a consequence of the smooth dependence of the function $G_{A}$ near $Q^{2}=0$, that can be linked with $P$-state contributions to the transitions form factors. This topic is discussed in Sec. V.4. The conclusion is then that near $Q^{2}=0$, at the scale of $Q^{2}=0.01$ GeV2 the axial-vector form factor $G_{A}$ has a slower falloff than the suggested by a dipole parametrization of $G_{A}$ with a typical cutoff $M_{A}\simeq 1.05$ GeV, for the $n\to p$ transition (0.13 fm2 to be compared with the experimental estimate 0.42 fm2).
The present results suggest then that $r_{A}^{2}\right>$ may decrease slightly in medium, contrary to what we naively observe, for instance by the calculations of the corresponding electromagnetic form factors, namely, charge radii of the octet baryons increase in medium [28, 29, 21, 27]. In contrast, the calculations of $r_{A}^{2}\right>$ from Refs. [81, 82] suggest that we may expect a small enhancement in-medium.
We notice, however, that the slow falloff of $G_{A}$ near $Q^{2}=0$ is a local effect, and when we extend the analysis to the range $Q^{2}=0.1$ GeV2, our model calculation for $G_{A}$ compares well with the dipole parametrization discussed above. Only for $Q^{2}>1.2$ GeV2 we observe significant differences from the dipole parametrization. Our model has a falloff slower than the dipole parametrization.
@@ V.3.2 Induced pseudoscalar form factor $G_{P}^{\ast}$
The study of the falloff of $G_{P}^{\ast}$ must be taken with care due the very rapidly changing function near $Q^{2}=0$, discussed in the previous sections. We make then a more qualitative analysis based on a ratio to a value of $Q^{2}$ in a range where the falloffs of $G_{P}$ are more well defined. We consider then the value $Q_{p}^{2}=0.012$ GeV2 (still a small value, large enough to avoid the turning effect of the function $G_{P}^{\rm B}$). The results of $G_{P}^{\ast}(Q^{2})/G_{P}^{\ast}(Q_{p}^{2})$ are presented on the each right side of Figs. 11 and 12 for $\Delta I=1$ and $\Delta S=1$, respectively.
The difference between the calculations for different $\rho$ for the case $\Delta I=1$ (right side of Fig. 11) are so difficult to observe, so we use a logarithmic scale. The similarity of the calculations is a consequence of the fact that the mass of the pion is unmodified by the medium and the factor $1/(m_{\pi}^{2}+Q^{2})$. From the figures, one can conclude that apart the value of the function $G_{P}$ near $Q^{2}=0$, the cases $\rho/\rho_{0}=0$, 0.5 and 1 are indistinguishable at low $Q^{2}$ ($Q^{2}<0.1$ GeV2). The ratio can then be interpreted as an universal function. The corollary of this result is that at low $Q^{2}$, we can determine the function $G_{P}^{\ast}(Q^{2})$ correcting the function in vacuum $G_{P}(Q^{2})$ by $G_{P}^{\ast}(Q_{p}^{2})/G_{P}(Q_{p}^{2})$.
@@
The results for $\Delta S=1$ (right side of Fig. 12), now in a linear scale, show a different behavior from $\Delta I=1$ cases, as a consequence of the factors $1/(m_{K}^{*2}+Q^{2})$. There are two typical cases, divided in the first two lines (light baryons) and the last two lines ($\Xi$ decays, heavy baryons) on the right side of Fig. 12. In the first case, for the $\Lambda\to p$ and $\Sigma^{-}\to n$/$\Sigma^{0}\to p$ transitions the ratios are almost the same near $Q^{2}=Q_{p}^{2}$, but differ in the range $Q^{2}\simeq 0.02$–0.2 GeV2, till the lines converge again. In that region the ratios are smaller for larger densities. In the second case, for the $\Xi^{-}\to\Lambda$ and $\Xi^{-}\to\Sigma^{0}$ transitions the ratios are the same below 0.5 GeV2, and differ after that point (larger ratio for larger densities). Differently from the case $\Delta I=1$ the different transitions are characterized by similar but different slopes for $G_{P}^{\ast}(Q^{2})/G_{P}(Q_{p}^{2})$ at low $Q^{2}$.
@@
The overall conclusion is that the functions $G_{P}^{\ast}(Q^{2})/G_{P}(Q_{p}^{2})$ have a very sharp variation at low $Q^{2}$, as a consequence of the significant effect of the pole terms associated with the pion and the kaon. In the case of $\Delta I=1$, we may expect that the ratio $G_{P}^{\ast}(Q^{2})/G_{P}(Q_{p}^{2})$ is the same for all transitions at low $Q^{2}$.
@@
The description of the function $G_{P}^{\ast}$ can be improved in the future, once stronger constrains are included in the determination of the component $G_{P}^{\rm B}$, and on the axial form factors in general.
V.4 Discussion

### v2

@@ V.2.2 Results for $G_{P}^{\ast}$
Examples of our calculations for the induced pseudoscalar form factor $G_{P}$ in the nuclear medium are presented on the right side of Fig. 5 for $|\Delta I|=1$ and Fig. 6 for $|\Delta S|=1$. The results for all transitions are presented in Appendix C (Figs. C1, C2 and C3). Examples of ratios $G_{P}^{\ast}/G_{P}$ are presented on the left side of Figs. 9 for $|\Delta I|=1$ and 10 for $|\Delta S|=1$. The comparison of the ratios for all transitions is presented in Appendix C (Figs. C4 and C5). The range of $Q^{2}$ is adjusted to each case in order to better visualize the shape of the functions.
@@
On the right side of Figs. 5, 6, C1, C2 and C3, we can still observe the enhancement of $G_{P}$ near $Q^{2}=0$ as in vacuum, for both $|\Delta I|=1$ and $|\Delta S|=1$. Since, as mentioned already, the $G_{P}$ estimates at $Q^{2}=0$ may be inaccurate, we avoid definitive conclusions about the point $Q^{2}=0$, and for the ratios $G_{P}^{\ast}/G_{P}$ in Figs. 9, 10, C4 and C5, we display only the region $Q^{2}>0.012$ GeV2.
@@
In general, one can observe an overall suppression of the form factors $G_{P}^{\ast}$ for both cases ($|\Delta I|=1$ and $|\Delta S|=1$) compared with the vacuum, particularly at low $Q^{2}$. The suppression increases with the density. The suppression at low $Q^{2}$ is diluted when $Q^{2}$ increases. In some cases, there are apparent overlaps of the different calculations ($\rho=0$, $0.5\rho_{0}$ and $\rho_{0}$), particularly for heavy baryons. See the last two lines on the right side of Fig. C1 and all lines on the right side of Fig. C3. The apparent overlaps are the consequence of the reduced scale in $Q^{2}$ and the thickness of the lines.
@@
The analysis of the suppression effect in medium on $G_{P}$ is simplified when we look at the ratios $G_{P}^{\ast}/G_{P}$ that are presented on the left side of Figs. 9 and C4 for $|\Delta I|=1$ and on the left side of Figs. 10 and C5 for $|\Delta S|=1$.
@@
In some cases ($\Sigma^{+}\to\Lambda$, $\Sigma^{0}\to\Sigma^{+}$, $\Xi^{-}\to\Lambda$ and $\Xi^{-}\to\Sigma^{+}$) we have more significant reductions of $G_{P}^{\ast}/G_{P}$ at low $Q^{2}$ due to the more effective reduction of the meson cloud component, as discussed already for $G_{A}^{\ast}$ (recall that $G_{P}^{\ast}\propto G_{A}^{\ast}$). At large $Q^{2}$, we can expect for sufficient large values of $Q^{2}$ the dominance of the pole term (proportional to $G_{A}^{{\rm B}\ast}$). This effect can be observed in the figures for the vacuum (left side of Figs. C1, C2 and C3) in the upper limit of $Q^{2}$.
@@
We start the discussion with the $|\Delta I|=1$ transitions. For the transitions $|\Delta I|=1$, on the left side of Figs. 9 and C4, the significant reduction of the ratio $G_{P}^{\ast}/G_{P}$ can be understood as a consequence of the dominance of the pole and meson cloud terms, a sum proportional to $G_{A}^{\ast}=G_{A}^{{\rm B}\ast}+G_{A}^{{\rm MC}\ast}$. In that case, $G_{P}^{\ast}\propto(M_{BB^{\prime}}^{\ast})^{2}G_{A}^{\ast}$, since in the present model $m_{\pi}$ has the same value for the vacuum and medium ($m_{\pi}$ is almost insensitive to the medium modifications [17, 18, 19, 20]).
@@
The result $G_{P}^{\ast}/G_{P}<1$ is then naturally explained by the mass reduction in medium ($\frac{M_{B}^{\ast}+M_{B^{\prime}}^{\ast}}{M_{B}+M_{B^{\prime}}}<1$) and by the reduction of $G_{A}$ in medium ($G_{A}^{\ast}/G_{A}<1$) discussed already. The suppression is more effective due to the power 2 in the mass ratio. One can then expect a triple suppression of the ratio $G_{P}^{\ast}/G_{P}$ in medium for $|\Delta I|=1$. The impact of the suppression is smaller for heavy baryons (mass ratio and $G_{A}^{\ast}/G_{A}$ are reduced more weakly), and lower densities $\rho$. Note, in this respect, that the reduction of $G_{P}^{\ast}/G_{P}$ is very mild for the $\Xi^{-}\to\Xi^{0}$ near the upper limit of $Q^{2}$.
We consider now the $|\Delta S|=1$ transitions. In this case, the pole and the meson cloud term include the factor $\frac{(M_{B}^{\ast}+M_{B^{\prime}}^{\ast})^{2}}{m_{K}^{*2}+Q^{2}}$. At low $Q^{2}$ the effect of mass reduction of the kaon cancels in part the effect of the reduction of the baryon masses in medium (factor $\frac{(M_{B}^{\ast}+M_{B^{\prime}}^{\ast})^{2}}{(M_{B}+M_{B^{\prime}})^{2}}$), since $\frac{m_{K}^{2}}{m_{K}^{\ast 2}}>1$. Thus, at low $Q^{2}$, the ratio is dominated by the reduction of the meson cloud contribution (more significant reduction at low $Q^{2}$). For large $Q^{2}$, the effect of the kaon mass modification $m_{K}\to m_{K}^{\ast}$ is included in the factor $1+\frac{m_{K}^{2}-m_{K}^{*2}}{m_{K}^{*2}+Q^{2}}$ and decreases quickly to the unit. The consequence is that ratio $G_{P}^{\ast}/G_{P}$ can be described by Eq. (38), with $G_{A}^{\ast}$ replaced by $G_{A}^{B\ast}$ (dominance of the pole term at large $Q^{2}$). The conclusion is then that at large $Q^{2}$, we can expect a stronger suppression for light baryons (more significant baryon mass effect and reduction of $G_{A}^{{\rm B}\ast}/G_{A}^{\rm B}$) and almost no medium effects for heavy baryons ($G_{P}^{\ast}/G_{P}\approx 1$).
@@ V.3 Axial form factors in medium – Falloffs with $Q^{2}$
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure 9: Induced pseudoscalar form factor $G_{P}$ for $|\Delta I|=1$ transitions. Ratios $G_{P}^{\ast}/G_{P}$ and $G_{P}^{\ast}(Q^{2})/G_{P}^{\ast}(Q_{p}^{2})$. The results for the $\Sigma^{+}\to\Lambda$ and $\Sigma^{0}\to\Sigma^{+}$ transitions are presented in Appendix C, Fig. C4.
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure 10: Induced pseudoscalar form factor $G_{P}$ for $|\Delta S|=1$ transitions. Ratios $G_{P}^{\ast}/G_{P}$ and $G_{P}^{\ast}(Q^{2})/G_{P}^{\ast}(Q_{p}^{2})$. The results for the $\Sigma^{-}\to n$, $\Sigma^{0}\to p$ and $\Xi^{-}\to\Lambda$ transition are presented in Appendix C, Fig. C5. The ratios are the same for $\Xi^{-}\to\Sigma^{0}$ and $\Xi^{0}\to\Sigma^{+}$.
@@ V.3.1 Axial-vector form factor $G_{A}^{\ast}$
Examples of calculations of $G_{A}^{\ast}(Q^{2})/G_{A}^{\ast}(0)$ are presented on the right side of Figs. 7 and 8, for $|\Delta I|=1$ and $|\Delta S|=1$, respectively. The results for all transitions are presented in Appendix B (Figs. B4 and B5).
@@
As mentioned already, the slope of $G_{A}^{\ast}(Q^{2})/G_{A}^{\ast}(0)$ is related to the axial-vector square radius $r_{A}^{2}\right>$ in the limit $Q^{2}=0$. The results displayed on Figs. 7, 8, B4, and B5 suggest, however, that the slope is smaller than that we can expect from a typical dipole form (2) with $M_{A}\simeq 1.05$ GeV: $r_{A}^{2}\right>=\frac{12}{M_{A}^{2}}\simeq 0.42$ fm2 for the $n\to p$ transition. The axial-vector square radius obtained for the vacuum is of the order of 0.1 fm2 for the $n\to p$, $\Sigma^{0}\to\Sigma^{+}$ transitions; 0.16 fm2 for $\Xi^{-}\to\Sigma^{0}$ and $\Xi^{0}\to\Sigma^{+}$ transitions; 0.2 fm2 for the $\Sigma^{+}\to\Lambda$ and about 0.35 fm2 for the $\Xi^{-}\to\Xi^{0}$, $\Lambda\to p$, $\Sigma^{-}\to n$ and $\Sigma^{0}\to p$. The $\Xi^{-}\to\Lambda$ transition is more atypical and has a negative axial-vector square radius (as can be inferred from the figure). These results are a consequence of the smooth dependence of the function $G_{A}$ near $Q^{2}=0$, that can be linked with $P$-state contributions to the transition form factors. This topic is discussed in Sec. V.5. The conclusion is then that near $Q^{2}=0$, at the scale of $Q^{2}=0.01$ GeV2 the axial-vector form factor $G_{A}$ has a slower falloff than suggested by a dipole parametrization of $G_{A}$ with a typical cutoff $M_{A}\simeq 1.05$ GeV, for the $n\to p$ transition (0.13 fm2 to be compared with the experimental estimate 0.42 fm2).
The present results suggest then that $r_{A}^{2}\right>$ may decrease slightly in medium, contrary to what we naively observe, for instance by the calculations of the corresponding electromagnetic form factors, namely, charge radii of the octet baryons increase in medium [28, 29, 21, 27]. In contrast, the calculations of $r_{A}^{2}\right>$ from Refs. [126, 127] suggest that we may expect a small enhancement in medium.
We notice, however, that the slow falloff of $G_{A}$ near $Q^{2}=0$ is a local effect, and when we extend the analysis to the range $Q^{2}=0.1$ GeV2, our model calculation for $G_{A}$ compares well with the dipole parametrization discussed above. Only for $Q^{2}>1.2$ GeV2, do we observe significant differences from the dipole parametrization. Our model has a falloff slower than the dipole parametrization.
@@ V.3.2 Induced pseudoscalar form factor $G_{P}^{\ast}$
The study of the falloff of $G_{P}^{\ast}$ must be taken with care due the very rapidly changing function near $Q^{2}=0$, discussed in the previous sections. We make then a more qualitative analysis based on a ratio to a value of $Q^{2}$ in a range where the falloffs of $G_{P}$ are more well defined. We consider then the value $Q_{p}^{2}=0.012$ GeV2 (still a small value, large enough to avoid the turning effect of the function $G_{P}^{\rm B}$). Examples of calculations of $G_{P}^{\ast}(Q^{2})/G_{P}^{\ast}(Q_{p}^{2})$ are presented on the right side of Fig. 9 for $|\Delta I|=1$, and on the right side of Figs. 10 for $|\Delta S|=1$. The calculations for all transitions are presented in Appendix C (see Figs. C4 and C5).
The difference between the calculations for different $\rho$ for the case $|\Delta I|=1$ (right side of Figs. 9 and C4) are difficult to observe, so we use a logarithmic scale. The similarity of the calculations is a consequence of the fact that the mass of the pion is unmodified by the medium and the factor $1/(m_{\pi}^{2}+Q^{2})$. From the figures, one can conclude that, apart from the value of the function $G_{P}$ near $Q^{2}=0$, the cases $\rho/\rho_{0}=0$, 0.5 and 1 are indistinguishable at low $Q^{2}$ ($Q^{2}<0.1$ GeV2). The ratio can then be interpreted as an universal function. The corollary of this result is that at low $Q^{2}$, we can determine the function $G_{P}^{\ast}(Q^{2})$ correcting the function in vacuum $G_{P}(Q^{2})$ by $G_{P}^{\ast}(Q_{p}^{2})/G_{P}(Q_{p}^{2})$.
@@
The results for $|\Delta S|=1$ (right side of Figs. 10 and C5), now in a linear scale, show a different behavior from $|\Delta I|=1$ cases, as a consequence of the factors $1/(m_{K}^{*2}+Q^{2})$. There are two typical cases, divided into light baryons ($\Lambda$ and $\Sigma$ decays) and into heavy baryons ($\Xi$ decays). In the first case, for the $\Lambda\to p$ and $\Sigma^{-}\to n$/$\Sigma^{0}\to p$ transitions, the ratios are almost the same near $Q^{2}=Q_{p}^{2}$ but differ in the range $Q^{2}\simeq 0.02$–0.2 GeV2, till the lines converge again. In that region the ratios are smaller for larger densities. In the second case, for the $\Xi^{-}\to\Lambda$ and $\Xi^{-}\to\Sigma^{0}$ transitions the ratios are the same below $Q^{2}=0.5$ GeV2, and differ after that point (larger ratio for larger densities). Differently from the case $|\Delta I|=1$, the different transitions are characterized by similar but different slopes for $G_{P}^{\ast}(Q^{2})/G_{P}(Q_{p}^{2})$ at low $Q^{2}$.
@@
The overall conclusion is that the functions $G_{P}^{\ast}(Q^{2})/G_{P}(Q_{p}^{2})$ have a very sharp variation at low $Q^{2}$, as a consequence of the significant effect of the pole terms associated with the pion and the kaon. In the case of $|\Delta I|=1$, we may expect that the ratio $G_{P}^{\ast}(Q^{2})/G_{P}(Q_{p}^{2})$ is the same for all transitions at low $Q^{2}$.
@@
The description of the function $G_{P}^{\ast}$ can be improved in the future, once stronger constrains are included in the determination of the component $G_{P}^{\rm B}$, and on the axial form factors, in general.
V.4 Summary of the results in medium
We can summarize now the numerical results for the weak interaction axial form factors, for the $|\Delta I|=1$ and $|\Delta S|=1$ transitions.
Axial-vector form factor $G_{A}^{\ast}$
From the discussion in Sec. V.2.1, we conclude that the axial-vector form factors are suppressed (quenched) in medium, since $G_{A}^{\ast}/G_{A}<1$ for the densities $\rho=0.5\rho_{0}$ and $\rho_{0}$. This suppression is a consequence of a reduction of the meson cloud contribution, more significant at low $Q^{2}$, and the bare contribution, more significant at large $Q^{2}$.
The suppression is stronger for light baryons (like in the $n\to p$ and $\Lambda\to p$ transitions) than for heavy baryons (like in the $\Xi^{-}\to\Xi^{0}$, $\Xi^{0}\to\Sigma^{+}$ and $\Xi^{-}\to\Sigma^{0}$ transitions) and increases with the density. At low $Q^{2}$, there is, in general, a stronger suppression of $G_{A}$ due to the suppression of the meson cloud contribution. In the case of light baryons, one can have a substantial suppression for intermediate $Q^{2}$ (about 30% for $n\to p$ and 25% for $\Lambda\to p$ for $Q^{2}=2$ GeV2 and $\rho=\rho_{0}$). In the case of the heavy baryons one has milder suppressions. For the transitions $\Xi^{-}\to\Xi^{0}$, $\Xi^{0}\to\Sigma^{+}$ and $\Xi^{-}\to\Sigma^{0}$ one can observe a suppression of 8%–12% at $\rho=\rho_{0}$, with a weak $Q^{2}$ dependence for the ratio $G_{A}^{\ast}/G_{A}$. For the intermediate mass cases (transitions $\Sigma^{+}\to\Lambda$, $\Sigma^{0}\to\Sigma^{+}$, $\Sigma^{-}\to n$, $\Sigma^{0}\to p$ and $\Xi^{-}\to\Lambda$) we observe intermediate suppressions.
Except for a slightly stronger suppression on the $|\Delta I|=1$ transitions, there are no significant differences in the magnitudes of $G_{A}^{\ast}/G_{A}$ between the $|\Delta I|=1$ and $|\Delta S|=1$ transitions.
Induced pseudoscalar form factor $G_{P}^{\ast}$
We summarize now the results from Sec. V.2.2 for the induced pseudoscalar form factor in medium. In general, one observes also a reduction of the ratio $G_{P}^{\ast}/G_{P}$ in the nuclear medium. In the present case, however, we need to distinguish between the results for the $|\Delta I|=1$ and $|\Delta S|=1$ transitions. The magnitude of the suppression increases with the density.
The difference between the $|\Delta I|=1$ and $|\Delta S|=1$ transitions is due to the dependence of the pole and meson cloud terms on the mass of the pion ($|\Delta I|=1$) or the mass of the kaon ($|\Delta S|=1$). As a consequence, $G_{P}$ has larger magnitudes near $Q^{2}=0$ for the $|\Delta I|=1$ transitions than for the $|\Delta S|=1$ transitions (the results for $|\Delta I|=1$ are about 10 times larger than the results for $|\Delta S|=1$). Another difference between the $|\Delta I|=1$ and $|\Delta S|=1$ transitions is that $G_{P}$ displays a much slower falloff with $Q^{2}$ for $|\Delta S|=1$, due to the factor $1/(m_{K}^{*2}+Q^{2})$ on the pole and meson cloud terms (the in-medium kaon mass is still larger than the pion mass).
In the case of $|\Delta I|=1$, the dominant term is the pole term. In these conditions, one can expect that $G_{P}^{\ast}/G_{P}\propto(M_{BB^{\prime}}^{\ast}/M_{BB^{\prime}})^{2}(G_{A}^{\ast}/G_{A})$. The consequence of this relation is that $G_{P}$ is suppressed in the nuclear medium and that the magnitude of the suppression is determined by the reduction of the masses, by the factor $(M_{BB^{\prime}}^{\ast}/M_{BB^{\prime}})^{2}$ and by the reduction of the $G_{A}^{\ast}$ in medium (factor $G_{A}^{\ast}/G_{A}$). The combination of the two effects is then a strong suppression for decays of light baryons and a not so strong suppression for decays of heavy baryons (smaller suppression factors).
As for the $|\Delta S|=1$ transitions, one observes also a reduction of $G_{P}$ in medium but with different trends for low and intermediate/large $Q^{2}$. At low $Q^{2}$, the magnitude of $G_{P}^{\ast}/G_{P}$ is smaller than the $|\Delta I|=1$ transitions, because in this case the reduced mass of the kaon tends to increase the magnitude of $G_{P}$ in-medium [$G_{P}^{\ast}/G_{P}\propto(m_{K}/m_{K}^{\ast})^{2}$], canceling partially the reduction due to the baryon masses and $G_{A}^{\ast}/G_{A}$. For larger values of $Q^{2}$, however, the variation of $m_{K}^{*}$ became irrelevant and we recover the relation $G_{P}^{\ast}/G_{P}\propto(M_{BB^{\prime}}^{\ast}/M_{BB^{\prime}})^{2}(G_{A}^{\ast}/G_{A})$ (valid for $|\Delta I|=1$). As a consequence, there is a significant suppression for light baryons and a very small suppression for heavy baryons (with almost no dependence on $Q^{2}$ at large $Q^{2}$).
In both cases ($|\Delta I|=1$ and $|\Delta S|=1$), the suppression on $G_{P}$ is stronger at low $Q^{2}$, due to the effect of the meson cloud (stronger suppression of $G_{A}$ in medium).
Combining the cases $|\Delta I|=1$ and $|\Delta S|=1$ one can have at intermediate $Q^{2}$ ($Q^{2}\approx 1$ GeV2) and $\rho=\rho_{0}$, reductions of about 30–40% for light baryons ($n\to p$, $\Lambda\to p$, $\Sigma^{-}\to n$ and $\Sigma^{0}\to p$) and about 5%–10% for heavy baryons ($\Xi^{-}\to\Xi^{0}$, $\Xi^{-}\to\Sigma^{0}$ and $\Xi^{0}\to\Sigma^{+}$). For intermediate mass baryons ($\Sigma^{+}\to\Lambda$, $\Sigma^{0}\to\Sigma^{+}$, $\Sigma^{-}\to n$, $\Sigma^{0}\to p$ and $\Xi^{-}\to\Lambda$) one has intermediate reductions (15%–20%).
V.5 Discussion